Find an equation for the hyperbola that satisfies the given conditions.

Foci: (0,+-3), hyperbola passes through (4, 5)
Answer #1

Solution-

Since foci lies along y - axis.So, hyperbola must be open along y - axis.

Let the equation for the hyperbola that satisfies the given conditions be

\frac{(y-k)^2}{a^2}+\frac{(x-h)^2}{b^2}=1.....(1)

Here, (h, k) = centre of hyperbola

a = distance between centre and vertex and

b = distance between centre and end point of conjugate axis.

Foci = (h, k ±c)

On comparing foci coordinates with given coordinates we get

Foci: = (h, k ±c) (0,±3)

h = 0 , k = 0 and c =3

On putting (h, k) = (0,0) in equation (1) to get

\frac{y^2}{a^2}-\frac{x^2}{b^2}=1 .....(2)

Now, Since c2 = a2 + b2 and c = 3

So, a2 + b2 = c2

Or a2 + b2 = 32

Or a2 + b2 = 9

Or b2 = 9 - a2 ......(3)

Since hyperbola passes through (4, 5)

So, putting (x,y) = (4,5) in equation (2) we get

\frac{5^2}{a^2}-\frac{4^2}{b^2}=1

\frac{25}{a^2}-\frac{16}{b^2}=1

Or 25b2 - 16a2 = a2b2

Or 25b2 - a2b2 -16a2 =0

Or b2( 25 - a2) - 16a2 =0

Putting the value of b2 from equation (3) to get

(9 - a2)(25 - a2) - 16a2 =0

Or 225 -9a2 -25a2 + a4 -16a2 =0

Or a4 -50a2 +225 =0

Solving it by factorisation method to get

a4 -45a2 -5a2 +225 = 0

a2(a2 -45) -5(a2 - 45) =0

(a2 - 45)(a2 - 5) =0

This implies

a2 = 45 , a2 = 5

Or a = √(45) , a= √5

Or a = 3√5 , a √5

Putting the values of a2 = 45, 5 in equation (3) to get

b2 = 9 - 45 = -36 (not possible) ,

b2 = 9 - 5 = 4 implies b = 2

So, a = √5 and b = 2

On putting a = √5 and b = 2 in equation (2) we get

The equation of hyperbola is

\frac{y^2}{\sqrt{5}^2}-\frac{x^2}{2^2}=1

\frac{y^2}{5}-\frac{x^2}{4}=1

Hence, the required equation of hyperbola is

\frac{y^2}{5}-\frac{x^2}{4}=1